OpenStudy (anonymous):

Sove this exponential equation to the nearest hundredth: 3^x+2=15

OpenStudy (rulnick):

Is the "+2" in the exponent?

OpenStudy (anonymous):

\[3^x=13\] \[x=\frac{\ln(13)}{\ln(3)}\]

OpenStudy (rulnick):

3^x=13 x = log13 / log3 = 2.33

OpenStudy (rulnick):

... and if the +2 is in the exponent it is 3^(x+2) = 15 (x+2) log 3 = log 15 x+2 = log15/log3 x = log15/log3-2 = 0.46